OpenStudy (anonymous):

Which fraction falls between five-fourteenths and five-sevenths? thirteen-fourteenths two-sevenths three-fourteenths four-sevenths

ganeshie8 (ganeshie8):

this one is a lil tricky, lets see..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have to go soon

OpenStudy (anonymous):

gtg

ganeshie8 (ganeshie8):

\[\dfrac{5}{14} < ?< \dfrac{5}{7}\] \[\dfrac{5}{14} < ?< \dfrac{10}{14}\]

OpenStudy (anonymous):

Thank you @ganeshie8 for all your help i have fanned you and i will medal you

ganeshie8 (ganeshie8):

np, feel free to tag me in ur future questions :)